Determine the curl of F which is (∇ × F ) at the point (2, 0, 3) given that



F = ze2xy i + 2xz cosy j+ (x + 2y)k

1
Expert's answer
2022-03-21T16:09:00-0400

"\text{curl}\vec F=(2-2x\cos y)\vec i+(e^{2xy}-1)\vec j+(2z\cos y-2xze^{2xy})\vec k,"

"\text{curl}\vec F(2,0,3)=-2\vec i-6\vec k."